### v2

@@ 3.3 Multimodal versatile attention mechanism
We leverage a multimodal versatile attention mechanism based on PyTorch FlexAttention [57] to empower OneCAT with the ability to process diverse modalities and tasks in a flexible and adaptive manner. As illustrated in Fig. 4, text tokens $T$ are processed using causal attention, ensuring autoregressive generation; Continuous visual tokens $U$ are processed via full attention, allowing each token to attend to all others in the sequence. Multiscale discrete visual tokens $G_{i}$ (where $i$ denotes the scale index) are processed via block causal attention, tokens within the same scale can attend to each other freely, while attention across scales follows a causal attention.
@@ 4 Model Training Pipeline
The training pipeline is divided into three stages and the instruction of the training recipe is shown in Tab. 1.
@@ 4.1.1 Stage 1-1: Teacher Training
Rather than employing an off-the-shelf MLLM as teacher (e.g., Qwen2.5-VL [3]), we construct a custom teacher model to ensure parameter consistency between the LLM backbones of the teacher and student models, thereby improving distillation efficiency. Specifically, the teacher is built by connecting a pre-trained vision encoder (InternViT [14]) and a LLM (Qwen2.5 [82]) with a two-layer MLP as connector. During this sub-stage, we freeze the ViT and LLM, and exclusively train the MLP connector on a curated small-scale dataset of image-to-text caption pairs using the standard NTP loss.
@@ 4.1.2 Stage 1-2: Expert Pretraining
With the teacher model prepared, we proceed to train the OneCAT model. We keep the QKV Projection, Attention, and Text FFN frozen, and selectively optimize the task-specific modules: the Und. FFN and Patch Embedding Layer for multimodal understanding, and the Gen. FFN for text-to-image generation.
@@
Understanding Distillation: We optimize the Und. FFN on a large-scale dataset of image-to-text caption pairs to acquire fundamental visual knowledge. Based on the specifically designed teacher, the training objective is a combination of the NTP loss ($\mathcal{L}_{\text{NTP}}$) and a distillation loss ($\mathcal{L}_{\text{Distill}}$). Specifically, $\mathcal{L}_{\text{NTP}}$ is the standard cross-entropy loss for autoregressive text generation. For distillation, instead of matching the final output logits, we align the student’s internal hidden states with those of the teacher model through deep feature-level matching. This strategy enables the student to not only mimic the teacher’s final prediction but also its intermediate computational patterns across all token positions (including both visual and text tokens) for better visual knowledge transfer. The distillation loss $\mathcal{L}_{\text{Distill}}$ is formulated as the sum of MSE losses between the hidden states of the student and teacher models over all $N$ transformer layers:
$\mathcal{L}_{\text{Distill}}=\sum_{n=1}^{N}\text{MSE}(\mathbf{h}_{S}^{(n)},\mathbf{h}_{T}^{(n)}),$ (1)
where $\mathbf{h}_{S}^{(n)}$ and $\mathbf{h}_{T}^{(n)}$ represent the hidden state outputs from the $n$-th transformer block of the student and teacher models, respectively. The final objective is thus:
$\mathcal{L}_{\text{Und}}=\mathcal{L}_{\text{NTP}}+\lambda\mathcal{L}_{\text{Distill}},$ (2)
where $\lambda=0.2$ is a balancing hyperparameter. Throughout distillation, all models process images at a fixed resolution of $448\times 448$ to balance computational load and the granularity of visual features.
Generation Pretraining: In parallel, we optimize the Gen. FFN on a delicate text-to-image generation dataset to enable the model to learn the spatial relationships and dependencies between multi-scale discrete visual tokens. We adopt the cross-entropy loss for next-scale prediction [66, 24] and the output image resolution is fixed to $\times256$ in this stage.
@@ 4.2 Stage-2: Unified Mid-Training
In the second stage, we unfreeze the entire model to achieve unified mid-training across multiple tasks (i.e., image-to-text, text-to-image, image editing, and text-only dialogues). At this stage, we incorporate the proposed scale-aware adapter (SAA), which is optimized with other modules of LLM together to extract scale-specific representation for enhanced image generation.
We introduce native resolution strategy for both understanding and generation in this stage. For multimodal understanding, the model is trained to process images at their original resolutions, thereby preserving fine-grained details and eliminating information loss. Additionally, a thumbnail of resolution $\times448$ is included to provide global visual context. For visual generation, the model is trained to generate images with different resolution and aspect ratios where the side lengths dynamically sampled from a range of 288 to 864 pixels, which significantly enhancing its generative versatility and real-world applicability.
@@ 4.3 Stage-3: Unified Supervised Fine-tuning
The final stage involves unified supervised fine-tuning (SFT) using a curated dateset of higher-quality data to enhance instruction-following and visual generation quality. The native resolution strategy was continued, with the size of generated image expanded to support side lengths between 288 to 1728 pixels, enabling higher-resolution results.
@@ 5 Data Setup
Stage-1: For the multimodal understanding, we curate a large-scale dataset of approximately 436 million image-text pairs, which is meticulously compiled and processed through comprehensive filtering and deduplication. This dataset is collected from two primary sources: (1) Public Available Image-Text Caption Pairs: We incorporate several publicly available, high-quality image-caption datasets, including Recap-DataComp-1B [40], Capsfusion [86], Detailed-Caption [41], SA1B-Dense-Caption [15], and Moondream2-COYO-5M-Captions [31]. (2) Re-captioned Image Datasets: We generate new image-text pairs by re-captioning large-scale public image collections. The source image datasets for this process include COYO700M [6], CC12M [7], CC3M [60], LAION-400M [59], and Zeor250M [78]. From this dataset, we randomly select a subset of 10 million samples to train the custom teacher.
@@
For image generation, we construct a dataset of 52 million text-to-image samples after a rigorous filtering process to remove samples with low resolution or poor aesthetic scores. This collection consists of 1 million class-labeled images from ImageNet-1k [17], 20 million pairs from re-captioned public collections (i.e., COYO700M [6], LAION-400M [59] and CC12M [7]), and 30 million in-house synthetic images generated by FLUX. The overall training token ratio across multimodal understanding and visual generation samples in Stage-I is approximately 8:1.
@@
Stage-2: In the unified mid-training, for multimodal understanding we leverage an in-house dataset of 70 million visual instruction samples. This dataset is specifically curated to be highly diverse tasks, including general VQA, detailed image captioning, OCR, multimodal reasoning(i.e., STEM problem-solving), knowledge, and visual grounding. For visual generation, we supplement the text-to-image samples of Stage-1 with a additional collection of 8 million image editing samples, resulting a total of 60 million visual generation samples. These additional image editing samples are sourced from several public image editing datasets, including AnyEdit [85], UltraEdit [93], HQ-Edit [30] and OmniEdit [72]. Additionally, we incorporate 40 million text-only instruction samples to preserve the language ability of LLM. To ensure a strong focus on visual generation in Stage-II, we oversample the visual generation data, resulting a final training token ratio of approximately 1 :2 :6 across text-only, multimodal understanding, and visual generation tasks, respectively.
OneCAT-1.5B OneCAT-3B
Base Model Qwen2.5-1.5B-instruct Qwen2.5-3B-instruct
Active Parameters 1.5B 3B
Total Parameters 4.5B 9B
Understanding Distillation
Teacher ViT InternViT-300M
Teacher LLM Qwen2.5-1.5B-instruct Qwen2.5-3B-instruct
@@
Table 2: Model configurations for the two variants of OneCAT.
@@
Stage-3: In the SFT stage, for multimodal understanding and text-only instruction, we construct a high-quality dataset of 13 million samples, combining 10 million from public MAmmoTH-VL dataset [22] with 3 million of our in-house synthetic instruction samples to further improve the reasoning abilities. For visual generation, we utilize a total of 3 million samples, aggregated from UniWorld [46], blip3o-60k [8], ShareGPT-4o-Image [10], and additional in-house synthetic data generated by GPT-4o and FLUX. The overall training token ratio across text-only, multimodal understanding, and visual generation is approximately 1 :5 :6.
@@ 6 Implementation details
Model Configurations: We conduct our experiments on two model variants, OneCAT-1.5B and OneCAT-3B. The OneCAT-1.5B model is based on Qwen2.5-1.5B-instruct [3] and contains 1.5B active parameters (4.5B total). The OneCAT-3B model is built upon Qwen2.5-3B-instruct [3] and utilizes 3B active parameters (9B total). For understanding distillation, the ViT of teacher model is InternViT-300M [14] for both two variants and the LLM of teacher model is Qwen2.5-1.5B-instruct and Qwen2.5-3B-instruct for OneCAT-1.5B and OneCAT-3B, respectively, to align with the LLM backbone of our OneCAT. Detailed instruction of the model configurations is shown in Tab. 2.
@@ 7 Evaluation
We evaluate our model on 12 public Multimodal understanding benchmarks spanning diverse capabilities: MMbench [49], MME [84], MMMU [88], MM-Vet [87], and SEED [36] assess general multimodal perception and reasoning. MathVista [50] focuses on mathematical reasoning. TextVQA [63], ChartQA [52], InfoVQA [54], and DocVQA [53] evaluate OCR and text-related visual question answering; while GQA [29] evaluates visual scene understanding and AI2D [33] evaluates scientific diagram comprehension.
@@
Figures 6 and 7 present qualitative comparisons for the text-to-image and image-editing tasks, respectively. OneCAT-3B exhibits leading instruction-following and world-understanding capabilities among open-source models. For example, in the fourth row of Figure 6, it is the only open-source model that correctly generates exactly four characters in a pixel-art style; in contrast, BAGEL-7B produces five characters, and Janus-Pro-7B fails to render pixel art. Similarly, in the last row of Figure 7, only OneCAT-3B produces an image with the correct lighting conditions.
In summary, the strong instruction-following capability in both text-to-image generation and image editing tasks highlights the effectiveness and versatility of our architectural design.
@@ 7.3 Comparison of Inference Efficiency
In addition to its strong performance, our model’s architectural design also yields significant improvements in inference efficiency for handling both high-resolution image input and output. We evaluated the efficiency of OneCAT-3B in two distinct phases: prefilling and generation.
In the prefilling phase, we measure the Time to the First Token (TTFT) to assess the computational cost of processing the input. As shown in Tab. 8, thanks to our pure decoder-only architecture that eliminates the need for a separate ViT encoder, OneCAT-3B demonstrates a substantial efficiency advantage, particularly when handling high-resolution images. Compared to QwenVL-3B, as the image resolution increases from $768\times 768$ to $1792\times 1792$, the reduction in TTFT for OneCAT-3B grows sharply from 50.4% to 61.4%. This provides strong evidence for the effectiveness of our design, especially for high-resolution inputs.
In the generation phase, we evaluated the total inference time for text-to-image (T2I) generation and image editing. As detailed in Tab. 9, OneCAT-3B’s inference speed far surpasses that of the strong diffusion-based baseline, BAGEL-7B. For instance, when generating a $1024\times 1024$ resolution image, OneCAT-3B’s T2I and editing inference times are merely 2.85s and 4.61s, respectively, making it approximately $10\times$ faster than BAGEL-7B. This significant advantage stems from our pioneering multi-scale autoregressive generation mechanism inside LLM that enables parallel token generation, which drastically reduces the number of required decoding steps compared to iterative diffusion models.
In summary, OneCAT achieves exceptional inference efficiency in both the prefilling and generation stages. This highlights the immense potential and practical value of our unified autoregressive framework for building efficient and powerful large multimodal models.
Table 8: Efficiency comparison of OneCAT-3B and Qwen2.5-VL-3B. Models are tested based on one NVIDIA H800 GPU. We report the time to the first token (TTFT) to measure the reduction of computational cost of prefilling phase. 256* denotes the number of visual tokens for thumbnail.
Model Resolution of Input Image #Input Text Tokens #Input Visual Tokens TTFT(s) Reduction
Qwen2.5-VL-3B $768\times 768$ 24 731 0.135 50.4%
OneCAT-3B $768\times 768$ 24 731 $256^{*}$ 0.067
Qwen2.5-VL-3B $1024\times 1024$ 24 1395 0.216 57.4%
OneCAT-3B $1024\times 1024$ 24 1395 $256^{*}$ 0.092
Qwen2.5-VL-3B $1792\times 1792$ 24 4098 0.583 61.4%
OneCAT-3B $1792\times 1792$ 24 4098 $256^{*}$ 0.225
Table 9: Generation efficiency comparison of OneCAT-3B and BAGEL. Models are tested based on one NVIDIA H800 GPU. We report total inference time to measure the Text-to-Image(T2I) and Image-Editing efficiency.
Model Resolution of Generated Image T2I Infer. Time (s) Edit Infer. Time (s)
@@
OneCAT-3B $1024\times 1024$ 2.85 4.61

### v3

@@ 3.3 Multimodal versatile attention mechanism
We employ a multimodal versatile attention mechanism based on FlexAttention [76] to enable flexible processing of diverse modalities and tasks within a single LLM. As shown in Fig. 4, text tokens $T$ use causal attention for autoregressive generation, while continuous visual tokens $U$ apply full attention for global interaction. Multiscale discrete visual tokens $G$ use block causal attention. Tokens within the same scale attend to each other freely, while attention across scales follows a causal attention.
@@ 4 Model Training Pipeline
As shown in Fig. 5 and Table 1, we employ a three-stage training pipeline for OneCAT.
@@ 4.1.1 Stage 1-1: Teacher Training
Instead of using an off-the-shelf MLLM as teacher (e.g., Qwen2.5-VL [5]), we customize a teacher to ensure parameter consistency between the LLM backbones of the teacher and student, and thus, improve distillation stability and efficiency (refer to Fig. 8). Specifically, the teacher is built by connecting a pre-trained ViT (InternViT [19]) and a LLM (Qwen2.5 [105]) with a two-layer MLP. We freeze both ViT and LLM and only train the MLP connector on a small-scale image-text caption dataset with NTP loss. This alignment training endows the custom MLLM teacher with strong visual perception ability.
@@ 4.1.2 Stage 1-2: Expert Pretraining
We then train OneCAT based on the teacher model. We freeze the QKV, Attention, and Text FFN, and optimize the task-specific modules: the Und. FFN and Patch Embedding layer for visual understanding, and the Gen. FFN for visual generation.
@@
Understanding Distillation: We optimize the Und. FFN on a large-scale dataset of image-to-text pairs. The training objective is a combination of the NTP loss ($\mathcal{L}_{\text{NTP}}$) and a distillation loss ($\mathcal{L}_{\text{Distill}}$). Specifically, $\mathcal{L}_{\text{NTP}}$ is the cross-entropy loss for text token generation. For distillation, instead of matching the final output logits, we align the student’s internal hidden states with those of the teacher model through deep feature-level matching across each transformer layer. This enables the student to not only mimic the teacher’s final text prediction but also its internal computational patterns across all token (including both visual and text tokens) for better visual knowledge transfer. The distillation loss is formulated as $\mathcal{L}_{\text{Distill}}=\sum_{n=1}^{N}\text{MSE}(\mathbf{h}_{S}^{(n)},\mathbf{h}_{T}^{(n)})$, where $\mathbf{h}_{S}^{(n)}$ and $\mathbf{h}_{T}^{(n)}$ represent the hidden state outputs from the $n$-th transformer block of the student and teacher models, respectively. The final objective is $\mathcal{L}_{\text{Und}}=\mathcal{L}_{\text{NTP}}+\lambda\mathcal{L}_{\text{Distill}}$, where $\lambda=0.02$ is a balancing hyperparameter. Throughout distillation, all models process images at a fixed resolution of $448\times 448$ to balance computational load and the granularity of visual features.
Generation Pretraining: In parallel, we optimize the Gen. FFN on a delicate text-to-image generation dataset to enable the LLM to learn the spatial relationships and cross-scale dependencies. We also adopt the cross-entropy loss for next-scale prediction [89] and the output image resolution is fixed to $\times256$.
@@ 4.2 Stage-2: Unified Mid-Training
In Stage-2, we unfreeze the entire model to perform unified mid-training across multiple tasks (i.e., image-to-text, text-to-image, image editing, and text-only dialogues). We incorporate the proposed scale-aware adapter in this stage, which is optimized with other modules together to extract scale-specific representation for enhanced image generation quality. We also introduce native resolution strategy in this stage. For visual understanding, the model is trained to process images at their original resolutions, thereby preserving fine-grained details and reducing information loss. Additionally, a thumbnail of resolution $\times448$ is included to provide global visual context. For visual generation, we train the model with dynamical resolution and aspect ratios where the side lengths sampled from a range of 288 to 864 pixels, which enhancing its generation versatility and real-world applicability.
@@ 4.3 Stage-3: Unified Supervised Fine-tuning
The final stage involves unified supervised fine-tuning (SFT) using a curated dateset of higher-quality data to enhance instruction-following and visual generation quality. The native resolution strategy was continued, with the size of generated image expanded to support side lengths between 288 to 1776 pixels, enabling higher-resolution results.
@@ 5 Data Setup
Stage-1: For the multimodal understanding, we curate a large-scale dataset of approximately 436 million image-text pairs, which is meticulously compiled and processed through comprehensive filtering and deduplication. This dataset is collected from two primary sources: (1) Public Available Image-Text Caption Pairs: We incorporate several publicly available, high-quality image-caption datasets, including Recap-DataComp-1B [52], Capsfusion [110], Detailed-Caption [53], SA1B-Dense-Caption [20], and Moondream2-COYO-5M-Captions [40]. (2) Re-captioned Image Datasets: We generate new image-text pairs by re-captioning large-scale public image collections using Qwen2-VL [92]. The source image datasets for this process include COYO700M [8], CC12M [9], CC3M [80], LAION-400M [79], and Zeor250M [101]. From this large-scale dataset, we randomly sample a small-scale subset of 10 million samples to train the custom teacher.
@@
For image generation, we construct a dataset of 52 million text-to-image samples after a rigorous filtering process to remove samples with low resolution or poor aesthetic scores. This collection consists of 1 million class-labeled images from ImageNet-1k [22], 20 million pairs from public collections (i.e., COYO700M [8], LAION-400M [79] and CC12M [9]), and 30 million in-house synthetic images generated by FLUX. The overall training token ratio across multimodal understanding and visual generation samples in Stage-I is approximately 8:1.
@@
Stage-2: In the unified mid-training, for multimodal understanding we leverage an curated dataset of 70 million visual instruction samples. This dataset is specifically curated to be highly diverse tasks, including general VQA, detailed image captioning, OCR, multimodal reasoning(i.e., STEM problem-solving), knowledge, and visual grounding, which are sourced from Detailed-Caption [53], ALLaVA [10], ShareGPT4V [14], SA1B-Dense-Caption [20], WIT [85], pdfa-eng-wds [1], UReader [106], DVQA [69], OCR-VQA [71], WebSRC [16], GQA [37], visual-genome [44], GRIT [74], and other in-house synthetic visual instruction data.
For visual generation, we supplement the text-to-image samples of Stage-1 with a additional collection of 8 million image editing samples, resulting a total of 60 million visual generation samples. These additional image editing samples are sourced from several public image editing datasets, including AnyEdit [109], UltraEdit [117], HQ-Edit [38] and OmniEdit [95].
Additionally, we incorporate 40 million text-only instruction samples to preserve the language ability of LLM. To ensure a strong focus on visual generation in Stage-II, we oversample the visual generation data, resulting a final training token ratio of approximately 1 :2 :6 across text-only, multimodal understanding, and visual generation tasks, respectively.
@@
Table 2: Model configurations for the two variants of OneCAT. A params indicates the activated parameters and T params indicates the total parameters.
Model Recipe Understanding Distillation
Base Model A Params T Params Teacher ViT Teacher LLM
OneCAT-1.5B Qwen2.5-1.5B-instruct 1.5B 4.5B InternViT-300M Qwen2.5-1.5B-instruct
OneCAT-3B Qwen2.5-3B-instruct 3B 9B InternViT-300M Qwen2.5-3B-instruct
Stage-3: In the SFT stage, for multimodal understanding and text-only instruction, we construct a high-quality dataset of 13 million samples. This dataset comprises 10 million filtered samples from MAmmoTH-VL dataset [29] and 3 million samples from other open-source datasets AI2D [43], OKVQA [66], VQAv2 [28], ART500K [65], ScienceQA [78], GQA [37], CLEVR-Math [59], COCO-ReM [84], TallyQA [2], Docmatix [46], DVQA [69], DreamSim [26], ShareGPT4o [18].
@@
For visual generation, we utilize a total of 3 million samples, aggregated from UniWorld [58], BLIP3o-60k [11], ShareGPT-4o-Image [13], and additional synthetic data generated by GPT-4o [39] and FLUX [45] using the partial prompts from JourneyDB [86]. The overall training token ratio across text-only, multimodal understanding, and visual generation for unified sft is approximately 1 :5 :6.
@@ 6 Implementation details
Model Configurations: We conduct experiments on two model variants, OneCAT-1.5B and OneCAT-3B, which is initialized from Qwen2.5-1.5B-instruct and Qwen2.5-3B-instruct, respectively. The total parameters count of OneCAT-3B is 9B, but the activated parameters count for each token is 3B during forward process. Refer to Table 2 for more details.
@@
Classifier-free guidance (CFG). We follow previous works [13, 21] to use CFG for enhanced visual generation quality. For training, we randomly drop conditional text and reference image tokens. For inference, we combines conditional and unconditional predicted logits. For more details, please refer to Appendix B.
@@ 7 Evaluation
We evaluate OneCAT on several multimodal understanding benchmarks: MMbench [62], MME [108], MMMU [112], MM-Vet [111], and SEED [48] assess general multimodal perception and reasoning. MathVista [63] focuses on mathematical reasoning. TextVQA [83], ChartQA [67], InfoVQA [70], and DocVQA [68] evaluate OCR and text-related visual question answering; while GQA [36] evaluates visual scene understanding and AI2D [42] evaluates scientific diagram comprehension.
@@
7.3 Qualitative results.
We present qualitative comparisons for the text-to-image and image-editing tasks in Fig. 6 and 7, respectively. OneCAT-3B exhibits leading instruction-following and world-understanding capabilities. We also present more visual generation and understanding showcases in Appendix E.
@@ 7.4 Comparison of Inference Efficiency
In addition to its strong performance, our model’s architectural design also yields significant improvements in inference efficiency for handling both high-resolution image input and output.
The left section of Tab. 8 compares the inference efficiency of OneCAT and Qwen2.5-VL for multimodal understanding. Benefiting from our pure decoder-only architecture, which removes the external ViT encoder, OneCAT achieves significantly faster prefilling, reducing first-token latency by up to 61% compared to Qwen2.5-VL on high-resolution inputs. As shown in the right section of Tab. 8, OneCAT exhibits a substantial speed advantage over the diffusion-based BAGEL in image generation. When producing a $1024\times 1024$ image, OneCAT requires only 2.85s for T2I generation and 4.61s for image editing—approximately $10\times$ faster than BAGEL. This improvement stems from two key elements: our multi-scale autoregressive mechanism within the LLM, and a VAE tokenizer-free design for editing that reduces both encoding and decoding overhead.
Table 8: Efficiency comparison for understanding (Left) and generation (Right), tested on one NVIDIA H800. Left: We report the time to first token (TTFT). The number of input text tokens are fixed to 24. 256* is the number of visual tokens of thumbnail. Right: We report total inference time for Text-to-Image (T2I) and Image-Editing.
Model Resolution of Input Image #Tokens Visual TTFT(s)
Qwen2.5-VL-3B $768\times 768$ 731 0.135
[…]
Qwen2.5-VL-3B $1792\times 1792$ 4098 0.583
[…]
Model Resolution of Output Image T2I Infer. Time (s) Edit Infer. Time (s)
@@
[…]
7.5 Ablation Study
7.5.1 Ablation Study on Understanding Distillation (Und. Distil.) Stage
We ablate each module in the Und. Distil. stage using the OneCAT-1.5B. Each experiment uses 8B sampled tokens ($\sim$10M image-text pairs from Stage-1 dataset) to optimize the Und. FFN, followed by a simplified SFT using LLaVA-665k [60] dataset.
On Different Distillation Strategies. As shown in Tab. 9 and Fig. 8(a), our proposed distillation approach—distilling all layers’ hidden states—significantly improves visual learning efficiency and multimodal performance. In addition, distilling only the last layer’s hidden states or logits also brings gains over the baseline, yet remains inferior to our full distillation approach. This highlights the necessity of full-layer distillation for effective visual knowledge transfer from the custom teacher MLLM. Notably, our strategy yields better performance than applying the distillation approaches of EvE [23] and VoRA [91] to our OneCAT-1.5B under the same setting, further validating the efficacy of our proposed method.
Methods MMB MME-S MMVet SEED AI2D ChartQA TextVQA Avg.
w/o distillation 42.0 1209 13.2 47.0 53.8 10.5 10.0 31.4
distill last layer’s logits 44.8 1312 15.4 53.0 55.2 11.5 10.6 33.9
distill last layer’s hidden states 46.0 1255 15.3 54.3 55.3 11.0 10.9 33.9
[…]
EvE [23] 44.5 1276 15.0 52.8 54.0 10.6 11.2 33.4
VoRA [91] 41.5 1234 15.1 50.6 54.2 10.4 10.5 32.3
Table 9: Effect of different distillation strategies.
Methods Avg.
w/o distillation 31.4
[…]
distill with Qwen2.5-VL teacher 33.7
Table 10: Effect of different teachers. Avg. denotes the average score in Tab. 9.
On Different Teachers. We then analyze the effect of different teacher models. As shown in Tab. 10 and Fig. 8(b), using the pretrained Qwen2.5-VL-1.5B as the teacher results in unstable training and lower distillation efficiency due to parameter misalignment between the teacher and student LLMs—especially for the frozen attention and QKV layers—highlighting the importance of our custom teacher design.
Figure 8: Comparision of different distillation strategies and teachers for stage-1 training.
7.5.2 Ablation Study on Unified Mid-Training (Stage-2)
We initialize the model with OneCAT-1.5B from Stage-1 and perform a simplified unified mid-training to study (i) the effect of token ratio across tasks and (ii) the effect of SAA. All models then undergo a simplified SFT using LLaVA-665k [60] for understanding and BLIP3o-60k [11] for generation with a 1:1 token ratio.
On Token Ratio. We sample 5B and 10B tokens for text-only (T) and multimodal understanding (U), and vary the number of trained tokens in visual generation (G) to study the impact of token ratio. As shown in Table 11, increasing the training tokens for G does not significantly affect U but improves G. To balance training cost and performance, we adopt a token ratio of 1:2:6 for Stage-2.
Token Ratio Understanding Generation
(T:U:G) MMB MME-S MMVet SEED AI2D ChartQA TextVQA Avg. GenEval DPG
5B : 10B : 0B 61.6 1556 28.9 67.2 60.1 56.1 56.2 55.1 - -
5B : 10B : 10B 62.7 1547 29.7 66.7 60.3 59.8 55.4 55.6 80.7 74.6
[…]
5B : 10B : 45B 62.1 1544 30.1 66.8 58.4 58.6 55.0 55.2 81.4 75.6
Table 11: Effect of trained token ratio across text-only (T), understanding (U), and generation (G).
GenEval DPG
[…]
w/o SAA 78.1 74.0
Table 12: Effect of SAA.
On SAA. We then remove the SAA during the stage-2 training for performance comparison, as shown in Table 12. Each model train 5B, 10B, 30B tokens for T, U, and G for Stage-2, respectively. We can see that removing SAA leads to an obvious performance drop. We also present the visualization of frequency properties of tokens from different scales in Appendix D to better understand the motivation of our SAA.